SkinnyMinnie20:

Use the net to find the surface area of the prism. 300 in2 584 in2 40 in2 96 in2

SkinnyMinnie20:

1 attachment
SkinnyMinnie20:

hold on before you help me

SkinnyMinnie20:

okay I tried to do it myself and kept getting the wrong answer

Vocaloid:

first we treat the middle part as a rectangle and do: 8*(5+12+13) then we add the triangles

Vocaloid:



Vocaloid:

area of triangle = (1/2) * 5 * 12 two triangles = 2 * ( 1/2) * 5 * 12 so your answer = 8*(5+12+13) + 2 * ( 1/2) * 5 * 12

SkinnyMinnie20:

okay I did the first step but I did not add the triangles

Vocaloid:

yeah, if it asks for surface area you must add everything if it asks for lateral area you can exclude the top/bottom sides

SkinnyMinnie20:

300

Vocaloid:

good!